OpenStudy (anonymous):

We are given the derivative of f: f′(x) = cos(x) + 10/x. Compute f(2) − f(1) Can someone please help??

OpenStudy (anonymous):

compute the integral of f′(x) = cos(x) + 10/x. after getting that f(x) do f(2) and f(1) then find f(2) − f(1)

OpenStudy (anonymous):

thanks Mark. It takes me a while to figure things out, so give me a few moments and I'll write if I have further questions.

OpenStudy (anonymous):

ok, YW

OpenStudy (anonymous):

have fun and enjoy :D

OpenStudy (anonymous):

@mark_o. so, the integral of f'(x) = cos(x) + 10/x is sin(x) = 10 ln(x) Correct? So, for f(2) we would write this as sin(2) + 10 ln(2) and f(1) = sin(1) = 10 ln(1)?? How am I doing so far??

OpenStudy (anonymous):

So, next which is a ridiculous question, how do I calculate f(2) - f(1)?

OpenStudy (anonymous):

hmm wait your calculator must be in radan mode, is it?

OpenStudy (anonymous):

heh. yes.

OpenStudy (anonymous):

I think I'm good now. THANK YOU!

OpenStudy (anonymous):

you got it :D

OpenStudy (anonymous):

YW have fun now and enjoy :D